OpenStudy (anonymous):

Given that a,b and c are possible real numbers,such that a+b+c = x,then which of the following is true? 1) (x-a)(x-b)(x-c) >= 8abc 2) 1/a + 1/b + 1/c >= 9/x 3) (x-a)(x-b)(x-c) >= 8/27a^3

myininaya (myininaya):

possible real numbers?

myininaya (myininaya):

is that right?

OpenStudy (anonymous):

yes a,b,c are possible real numbers

myininaya (myininaya):

im just confused why it would say possible instead of saying let a,b,c be complex numbers

OpenStudy (anonymous):

a,b and c can be any real number its not a complex number for sure.

OpenStudy (anonymous):

Yeah, it sounds like it means arbitrary real number, not that the number might be real.

myininaya (myininaya):

well its possible for a,b, and c to be real but there is also the possibility of it being complex we arent sure what a b and c are

myininaya (myininaya):

so it doesn't say positive right?

OpenStudy (anonymous):

yes there can be possibility of a,b,c being complex numbers.but the question ask for the solution considering a,b,c being real and positive only

myininaya (myininaya):

i don't see positive anywhere i see possible which means we aren't sure what a b and c are

myininaya (myininaya):

so you are saying we get to assume a, b, and c are postive real numbers

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

You would have to assume that otherwise you get contradictions.

OpenStudy (anonymous):

can you please let me know how you came to conlusion

OpenStudy (anonymous):

how did you find the solution.???

OpenStudy (anonymous):

The first one does seem to work out.

OpenStudy (anonymous):

consider a=1,b=2,c=3 => x=6 ; option 1 and 2 both will get satisfied

OpenStudy (anonymous):

Yes, the second one seems to work as well. Both work for positive and negative test I've tried. Haven't tried the third, but there must be a more rigorous way of solving this.

OpenStudy (anonymous):

yup...!

OpenStudy (somethingawesome):

What you want to do is eliminate the incorrect solutions by coming up with a,b,c that makes the inequality fail. For the third one, a = 1, b = 0, c= 0 makes x = 1, so the left side is 0, the other is 8/27, so that one fails. For the second one, try a =1, b = -1, c = 2. Then 1/a+1/b+1/c = 1/2, but 9/x = 9/2, which is bigger, so that fails too. This leaves only the first one as a possibility.

myininaya (myininaya):

(x-a)(x-b)(x-c) =(x^2-ax-bx+ab)(x-c) =x^3-ax^2-bx^2+abx-x^2x+axc+bxc-abc =x^3-x^2(a+b+c)+x(ab+ac+bc)-abc =x^3-x^2*x+x(ab+ac+bc)-abc =x^3-x^3+x(ab+ac+bc)-abc =0+(a+b+c)(ab+ac+bc)-abc =(a+b+c)(ab+ac+bc)-abc =a^2b+a^2c+abc+ab^2+abc+b^2c+abc+ac^2+bc^2    -abc =2abc+a^2(b+c)+b^2(a+c)+c^2(a+b) =2abc+a(b^2+c^2)+b(a^2+c^2)+c(a^2+b^2) >2abc+a(b+c)^2+b(a+c)^2+c(a+b)^2 =2abc+a(b^2+2cb+c^2)+b(a^2+2ac+c^2)+ c(a^2+2ab+b^2) =8abc+a(b^2+c^2)+b(a^2+c^2)+c(a^2+b^2) >8abc assuming a>0,b>0,c>0

OpenStudy (anonymous):

that a great way of solving..!like it

myininaya (myininaya):

<2abc+a(b+c)^2+b(a+c)^2+c(a+b)^2* i made a mistake up there so this changes things :(

OpenStudy (anonymous):

none of the option will get satisied if you will take a=b=c=1